# Processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in Inert Higgs Doublet Models and Two Higgs Doublet Models

### v2

@@ Abstract
In this paper, we present the results for one-loop induced processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ with CP-even Higgses $\phi_{i,j}\equiv h,\leavevmode\nobreak\ H$ at high energy photon-photon collision, within the frameworks of Inert Higgs Doublet Models and Two Higgs Doublet Models. Total cross-sections are shown as functions of center-of-mass energies. We find that the cross-sections for the computed processes in all the models under investigations are enhanced at around the threshold of singly charged Higgs pair ($\sim 2M_{H^{\pm}}$). Furthermore, the enhancement factors defined as the ratio of cross-sections of $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the investigated models over the corresponding ones for $\gamma\gamma\rightarrow hh$ in the Standard Model, are examined in the model’s parameter space. In the Inert Higgs Doublet Models, the factors are studied in the parameter space of $(M_{H^{\pm}},\leavevmode\nobreak\ \mu^{2}_{2})$ and $(M_{H^{\pm}},\leavevmode\nobreak\ \lambda_{2})$. In the Two Higgs Doublet Models, the factors are examined in the planes of $(M_{H^{\pm}},\leavevmode\nobreak\ t_{\beta})$ as well as in the space of charged Higgs mass $M_{H^{\pm}}$ and the soft-breaking $Z_{2}$ parameter $m_{12}^{2}$. Two scenarios of $c_{\beta-\alpha}>0$ and $c_{\beta-\alpha}<0$ are studied in further detail. The factors give a different behavior from considering these scenarios. As a result, discriminations for the above-mentioned scenarios can be performed at future colliders.
@@ 1 Introduction
Measuring for scalar Higgs self-couplings including Standard Model-like Higgs trilinear- and quadratic-couplings as well as the couplings between scalar Higgses in many physics beyond the Standard Models (BSM) plays a key role for determining the scalar potential. We can subsequently answer for the origin of the electroweak symmetry breaking (EWSB). In this scheme, Higgs boson pair productions and multi-scalar Higgs productions should be measured precisely at future colliders. Recently, search for Higgs boson pair productions in the two bottom quarks associated with two photons, four bottom quarks, etc in the final states in proton-proton collisions have been performed at the Large Hadron Colliders (LHC) as in [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. It is well-known that the measurements for Higgs self-couplings are rather challenging at the LHC. The results from the study of [13, 14] show that the expected accuracy in the measurement of trilinear Higgs self-couplings would be about $20\%-30\%$ at the high luminosity of $3000$ fb-1. We know that physics the future lepton colliders (LC) will be complementary to the LHC in many aspects, as studied in [15]. Furthermore, the LC can significantly improve the LHC measurements in many cases and more important photon-photon collision is an option of the LC [16, 17] which the scalar Higgs pair productions ($\phi_{i}\phi_{j}$) can be measured via the channels $f\bar{f}\rightarrow f\bar{f}\gamma^{*}\gamma^{*}\rightarrow f\bar{f}\phi_{i}\phi_{j}$ for $f\equiv e,\mu$. In this aspect, the LC could open a window for probing new physics signals through multi-scalar Higgs productions.
@@
From theoretical calculation side, one-loop corrections to double Higgs productions at the LHC have calculated in Standard Model (SM), the Higgs Extensions of the Standard Models (HESM), as well as other BSM frameworks by many groups. For examples, it is worth to mention typical works as in [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78], for further reviews in [79] and the related-references therein. Additionally, one-loop induced for $hh$ productions in the high-energy $\gamma\gamma$ collisions in the SM, the HESM and other BSMs have computed in Refs. [80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94]. Other computations for one-loop induced for Higgs boson pair productions in the linear lepton colliders including future multi-TeV muon colliders, have performed in Refs. [95, 96, 97, 98] and the additional references therein. Furthermore, double pseudo-scalar Higgs $A^{0}A^{0}$ at a $\gamma\gamma$ collider in the Two Higgs Doublet Model have reported in Ref. [99].
In this paper, we present the results for one-loop induced processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ with CP-even Higgses $\phi_{i,j}\equiv h,\leavevmode\nobreak\ H$ at high-energy $\gamma\gamma$ collision within the Higgs Extensions of the Standard Models (HESM) including the Inert Higgs Doublet Models the Two Higgs Doublet Models. A general analytic formulas for the process amplitudes derived in ’t Hooft-Feynman gauge (HF) are valid for a class of the above-mentioned HESM. Analytic results for the calculated processes are presented via the scalar Passarino-Veltman (PV) functions following the output of the packages LoopTools [100] and Collier [101]. The analytical expressions are tested by several numerical checks , e.g the ultraviolet finiteness, infrared finiteness of the one-loop amplitudes. Furthermore, the amplitudes also obey the ward identity which the indentity is also verified numerically in the works. Additionally, both the packages LoopTools and Collier are used for cross-checking for the final results before generating physical results. In phenomenological results, cross-sections are shown as functions of center-of-mass energies. Furthermore, the enhancement factors defined as the ratio of cross-sections of $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the HESMs over the corresponding ones for $\gamma\gamma\rightarrow hh$ in the Standard Model, are examined in parameter space of the models under consideration.
The paper is structured as follows. We review briefly the HESM in the next section. We then present calculations for $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ with CP-even Higgses $\phi_{i,j}\equiv h,H$ in the HESM in the section $3$. The phenomenological studies for the HESM are discussed in section $4$. In section $5$, conclusion and outlook for the paper are shown. In appendices $A,\leavevmode\nobreak\ B$ we derive the couplings appear in the calculations.
@@ 2 Higgs Extensions of the Standard Model
Two typical Higgs Extensions of the Standard Models are studied in this paper. The first model is to the Inert Higgs Doublet Models which are reviewed in next subsection $2.1$. We then discuss the Two Higgs Doublet Models in subsection $2.2$.
@@ 2.1 Inert Higgs Doublet Models
In the IHDM, an inert scalar $SU(2)_{L}$ doublet is included into the potential of the SM. The inert scalar particles will respond for dark matter candidates. For reviewing the theory and phenomenological examinations for the IHDM in concrete, we cite to the following papers [104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116]. The scalar potential of the IHDM read the general expression as follows:
@@
The potential is conserved with respect to the so-called the global $Z_{2}$-symmetry, e.g. $\phi_{1}\leftrightarrow+\phi_{1}$, $\phi_{2}\leftrightarrow-\phi_{2}$. In this case, the scalar $\phi_{2}$ is odd and $\phi_{1}$, all particles in the SM are even under the $Z_{2}$-transformation. As mentioned, the $Z_{2}$-symmetry is unbroken after the EWSB, the field $\phi_{2}$ has the zero of vacuum expectation values (VEV). Otherwise, the field $\phi_{1}$ develops non-zero VEV ($v$). Two scalar fields are then expanded around VEV for the EWSB as follows:
@@
Where the VEV is fixed at $v\sim 246$ GeV (as the SM case). The Goldstone bosons $G^{\pm},\;G^{0}$ are giving the masses for $W^{\pm}$ boson and $Z$ boson, respectively. There is no mixing between $h$ and $H$. The physical spectrum of the IHDM after the EWSB is consisted of three neutral scalar physical states, e.g two CP-even Higgses $h,\leavevmode\nobreak\ H$ and a CP-odd Higgs $A_{0}$. In further, we have pair of singly charged Higgs bosons $H^{\pm}$ in this model. In the spectrum, neutral scalar Higgs $h$ is identical to the SM-like Higgs boson discovered at the LHC. The masses of scalar bosons are calculated from the pare parameters as follows:
@@
Where we have used $\lambda_{L/R}=\lambda_{3}+\lambda_{4}\pm\lambda_{5}$. As we mentioned in the above paragraphs, the global $Z_{2}$-symmetry is unbroken after the EWSB, The "inert" Higgs bosons like $H^{\pm}$, $A^{0}$ and $H$ have odd number under the $Z_{2}$-transformation. Subsequently, the "inert" Higgs bosons don’t couple to the SM particles. Therefore, the lightest neutral scalar bosons may be considered as dark matter candidates.
As a consequent of the unbroken of the $Z_{2}$-symmetry, the Yukawa Lagrangian of the IHDM must be the same as that of the SM. In detail, the Yukawa Lagrangian is expressed as follows:
@@
where the Yukawa coupling is given $g_{hff}=m_{f}/v$ for fermion $f$. All the couplings involving to the computed processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the IHDM are listed in Table 1 (for all physical couplings) and Table 2 (for unphysical particles). We emphasize that the processes $\gamma\gamma\rightarrow hH$ are forbidden in the IHDM due to the $Z_{2}$-symmetry. Therefore, we have only $\phi_{i}\phi_{j}\equiv hh,\leavevmode\nobreak\ HH$ in this case. The detailed expressions for all the concerned-couplings are derived in the appendix $A$.
@@
We are going to review the current constraints on the physical parameter space in the IHDM given in Eq. 13. The constraints for the physical parameter space can be obtained by including the theoretical conditions and the experimental data. In the perspective of the experimental data, we take into account the Electroweak Precision Tests (EWPT) of the IHDM, dark matter search at the LHC, as well as including the LEP data. The topics have studied in Refs. [104, 105, 106, 107]. Additionally, the implications for loop-induced decays of the SM-like Higgs ($h$) to $V\gamma$ with $V\equiv Z,\gamma$ in the IHDM framework, e.g. decay process $h\rightarrow\gamma\gamma$ have reported in [108, 115, 116], and decay chanels $h\rightarrow Z\gamma$ have examined in [115, 116]. Furthermore, searching signals of the IHDM at future colliders have performed in Refs. [109, 110, 111, 112, 113, 114]. In the theoretical side, the most important theoretical constraints are obtained from the conditions that the models follow the tree-level unitarity, the vacuum stability, the perturbative regime. The theoretical constraints give the limitations on the Higgs self-couplings $\lambda_{i}$ for $i=1,2,\cdots,5$ and $\mu_{2}$. Taking theoretical and experimental constraints in the above-mentioned papers, one can select the parameter space for the IHDM as follows: we can take $5$ GeV $\leq M_{H}\leq 150$ GeV, $70$ GeV $\leq M_{H^{\pm}},M_{A^{0}}\leq 1000$ GeV, $|\mu_{2}|\leq 500$ GeV, and $0\leq\lambda_{2}\leq 8\pi$.
@@ 2.2 Two Higgs Doublet Models
We next to consider the second kind of the Higgs extension of the SM which is to the Two Higgs Doublet Models (THDM) in this paper. For reviewing the theory and the phenomenological studies for the THDM, we refer the paper Ref. [117] for further information. The model is summarized briefly in this section. An complex Higgs doublet possessing hypercharge $Y=1/2$ is added into the scalar sector of the SM. The scalar potential is read the form of
@@
In the present work, the CP-conserving case of the THDM is examined. Subsequently, the pare parameters in the above potential are set to be real parameters in this version. Additionally, the potential of the THDM follows the $Z_{2}$-symmetry, e.g. $\phi_{1}\leftrightarrow\phi_{1}$ and $\phi_{2}\leftrightarrow-\phi_{2}$, up to the soft breaking terms as $m_{12}^{2}\phi_{1}^{\dagger}\phi_{2}+{\rm h.c.}$ are added into the potential. Where the parameter $m_{12}^{2}$ is the breaking scale for the $Z_{2}$-symmetry.
Two scalar doublets are exanded around their VEVs for the EWSB as
@@
The vacuum expectation value is then fixed at $v=\sqrt{v_{1}^{2}+v_{2}^{2}}\sim 246$ GeV in agreement with the SM case. After the EWSB, the physical particles in the THDM include of two CP-even Higgs bosons $h$ and $H$ in which one of them $h$ being SM-like Higgs boson found at LHC, a CP-odd Higgs ($A^{0}$) boson, and a pair of charged Higgses ($H^{\pm}$). To obtain the physical masses for the new scalar bosons, we perform the following rotations
@@
The mixing angle $\beta$ is given by $t_{\beta}\equiv\tan\beta=v_{2}/v_{1}$. The physical masses of Higgs bosons are then presented via the pare parameters as follows:
@@
We show the couplings concerning in the amplitude computations for the processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in Tables 34 (for physical couplings) and in Table 5 (for unphysical couplings). The couplings are derived in the appendix $B$.

### v3

@@ Abstract
In this paper, we present a phenomenological analysis of one-loop induced processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$, where the CP-even Higgs bosons are denoted as $\phi_{i,j}\equiv h,\leavevmode\nobreak\ H$, in high-energy photon-photon collisions, within the frameworks of the Inert Higgs Doublet Model and the Two Higgs Doublet Model. The total cross sections are evaluated as functions of the center-of-mass energy, finding that the cross sections for the considered processes in all the models under investigation are significantly enhanced around the threshold for charged Higgs boson pair production ($\sim 2M_{H^{\pm}}$). Furthermore, the enhancement factors, defined as the ratios of cross sections of $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the investigated models to those for $\gamma\gamma\rightarrow hh$ in the Standard Model, are examined in the relevant regions of the model’s parameter space. In the Inert Higgs Doublet Model, the factors are studied in the parameter space of $(M_{H^{\pm}},\leavevmode\nobreak\ \mu^{2}_{2})$ and $(M_{H^{\pm}},\leavevmode\nobreak\ \lambda_{2})$. In the Two Higgs Doublet Model, the factors are examined in the planes defined by $(M_{H^{\pm}},\leavevmode\nobreak\ t_{\beta})$ as well as in the space of the charged Higgs mass $M_{H^{\pm}}$ and the soft-breaking $Z_{2}$ parameter $m_{12}^{2}$. Two scenarios characterized by $c_{\beta-\alpha}>0$ and $c_{\beta-\alpha}<0$ are studied in further detail. The factors exhibit distinct behaviors between these two scenarios. As a result, it is possible to discriminate between them at future colliders. The dependence of the cross section for the process $\gamma\gamma\rightarrow hH$ on $m_{12}^{2}$ provides a potential probe of the soft $Z_{2}$-breaking scale in the Two Higgs Doublet Model.
@@ 1 Introduction
Measuring scalar Higgs self-couplings, including Standard Model-like Higgs trilinear and quadratic couplings, as well as the couplings between scalar Higgses in many beyond-the-Standard-Model (BSM) scenarios, plays a key role in determining the scalar potential. We can subsequently answer the question of the origin of electroweak symmetry breaking (EWSB). In this scheme, Higgs boson pair production and multi-scalar Higgs production should be measured precisely at future colliders. Recently, searches for Higgs boson pair production in final states such as two bottom quarks associated with two photons, four bottom quarks, etc., in proton-proton collisions have been performed at the Large Hadron Collider (LHC), as in [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. It is well known that measurements of Higgs self-couplings are rather challenging at the LHC. The results from the study in [13, 14] show that the expected accuracy in the measurement of trilinear Higgs self-couplings would be about $20\%$$30\%$ at the high luminosity of $3000$ fb-1. We know that the physics program of future lepton colliders (LC) will be complementary to that of the LHC in many aspects, as studied in [15]. Furthermore, the LC can significantly improve LHC measurements in many cases. More importantly, photon-photon collisions are an option at the LC [16, 17], where scalar Higgs pair production ($\phi_{i}\phi_{j}$) can be measured via the channels $f\bar{f}\rightarrow f\bar{f}\gamma^{*}\gamma^{*}\rightarrow f\bar{f}\phi_{i}\phi_{j}$ for $f\equiv e,\mu$. In this regard, the LC could open a window for probing new physics signals through multi-scalar Higgs production.
@@
From a theoretical calculation perspective, one-loop corrections to double Higgs production at the LHC have been calculated in the Standard Model (SM), the Higgs Extensions of the Standard Models (HESM), as well as in other BSM frameworks by many research groups. For example, it is worth mentioning notable works in [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78], for further reviews in [79] and the related references therein. Additionally, one-loop corrections for $hh$ production in high-energy $\gamma\gamma$ collisions in the SM, HESM, and other BSM models have been computed in Refs. [80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94]. Other computations for one-loop corrections to Higgs boson pair production at future linear lepton colliders, including multi-TeV muon colliders, have been performed in Refs. [95, 96, 97, 98] and the additional references therein. Furthermore, double pseudo-scalar Higgs $A^{0}A^{0}$ production at a $\gamma\gamma$ collider in the Two Higgs Doublet Model has been reported in Ref. [99]. One-loop contributions to the process $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ with CP-even Higgses $\phi_{i,j}\equiv h,\leavevmode\nobreak\ H$ in high-energy $\gamma\gamma$ collisions, valid for the Higgs Extensions of the Standard Models (HESM), are necessary. In our previous paper, analytic formulas for the concerned processes were reported [100].
In this paper, we present the phenomenological analysis of these processes at high-energy $\gamma\gamma$ collisions in the Inert Higgs Doublet Model (IHDM) and the Two Higgs Doublet Model (THDM). In the phenomenological analysis, cross sections are shown as functions of center-of-mass energies. Furthermore, the enhancement factors, defined as the ratios of cross sections of $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the HESMs to the corresponding ones for $\gamma\gamma\rightarrow hh$ in the Standard Model, are examined in the parameter space of the models under consideration. The paper is structured as follows. We briefly review the HESM in the next section. The phenomenological studies for the HESM are discussed in section 3. In section 4, we present the conclusion and outlook of the paper. In appendices A and B, we derive the couplings that appear in the calculations.
@@ 2 Higgs Extensions of the Standard Model
Two typical Higgs extensions of the Standard Model are studied in this paper. The first model is the Inert Higgs Doublet Model, which is reviewed in subsection 2.1. We then discuss the Two Higgs Doublet Model in subsection 2.2.
@@ 2.1 Inert Higgs Doublet Models
In the IHDM, an inert scalar $SU(2)_{L}$ doublet is included in the potential of the SM. The inert scalar particles serve as dark matter candidates. For a review of the theory and phenomenological studies of the IHDM in detail, we refer to the following papers [101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113]. The scalar potential of the IHDM is expressed as follows:
@@
The potential is conserved with respect to the so-called global $Z_{2}$-symmetry, e.g., $\phi_{1}\leftrightarrow+\phi_{1}$, $\phi_{2}\leftrightarrow-\phi_{2}$. In this case, the scalar $\phi_{2}$ is odd, and $\phi_{1}$ and all particles in the SM are even under the $Z_{2}$-transformation. As mentioned, the $Z_{2}$-symmetry remains unbroken after the EWSB. The field $\phi_{2}$ then has a zero vacuum expectation value (VEV), while the field $\phi_{1}$ develops a non-zero VEV ($v$). The two scalar fields are then expanded around their VEVs for the EWSB as follows:
@@
Where the VEV is fixed at $v\sim 246$ GeV (as in the SM case). The Goldstone bosons $G^{\pm},\;G^{0}$ give the masses for the $W^{\pm}$ boson and the $Z$ boson, respectively. There is no mixing between $h$ and $H$. The physical spectrum of the IHDM after the EWSB consists of three neutral scalar physical states: two CP-even Higgses $h,\leavevmode\nobreak\ H$ and a CP-odd Higgs $A_{0}$. Furthermore, we have a pair of singly charged Higgs bosons $H^{\pm}$ in this model. In the spectrum, the neutral scalar Higgs $h$ is identical to the SM-like Higgs boson discovered at the LHC. The masses of the scalar bosons are calculated from the model parameters as follows:
@@
We have used $\lambda_{L/R}=\lambda_{3}+\lambda_{4}\pm\lambda_{5}$. As mentioned in the previous paragraphs, the global $Z_{2}$-symmetry is unbroken after the EWSB. The "inert" Higgs bosons such as $H^{\pm}$, $A^{0}$, and $H$ have an odd number under the $Z_{2}$-transformation. Subsequently, the "inert" Higgs bosons do not couple to the SM particles. Therefore, the lightest neutral scalar bosons may be considered as dark matter candidates.
As a consequence of the unbroken $Z_{2}$-symmetry, the Yukawa Lagrangian of the IHDM must be the same as that of the SM. In detail, the Yukawa Lagrangian is expressed as follows:
@@
where the Yukawa coupling is given by $g_{hff}=m_{f}/v$ for fermion $f$. All the couplings involving the computed processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the IHDM are listed in Table 1 (for all physical couplings) and Table 2 (for unphysical particles). We emphasize that the processes $\gamma\gamma\rightarrow hH$ are forbidden in the IHDM due to the $Z_{2}$-symmetry. Therefore, we have only $\phi_{i}\phi_{j}\equiv hh,\leavevmode\nobreak\ HH$ in this case. The detailed expressions for all the concerned couplings are derived in appendix A.
@@
We are going to review the current constraints on the physical parameter space in the IHDM given in Eq. 13. The constraints for the physical parameter space can be obtained by including the theoretical conditions and the experimental data. In the perspective of the experimental data, we take into account the Electroweak Precision Tests (EWPT) of the IHDM, dark matter search at the LHC, as well as the LEP data. The topics have been studied in Refs. [101, 102, 103, 104]. Additionally, the implications for loop-induced decays of the SM-like Higgs ($h$) to $V\gamma$ with $V\equiv Z,\gamma$ in the IHDM framework, e.g. the decay process $h\rightarrow\gamma\gamma$, have been reported in [105, 112, 113], and decay channels $h\rightarrow Z\gamma$ have been examined in [112, 113]. Furthermore, searching for signals of the IHDM at future colliders has been performed in Refs. [106, 107, 108, 109, 110, 111]. In the theoretical side, the most important theoretical constraints are obtained from the conditions that the models follow tree-level unitarity, vacuum stability, and the perturbative regime. The theoretical constraints give the limitations on the Higgs self-couplings $\lambda_{i}$ for $i=1,2,\cdots,5$ and $\mu_{2}$. Taking the theoretical and experimental constraints from the above-mentioned papers, one can select the parameter space for the IHDM as follows: we can take $5$ GeV $\leq M_{H}\leq 150$ GeV, $70$ GeV $\leq M_{H^{\pm}},M_{A^{0}}\leq 1000$ GeV, $|\mu_{2}|\leq 500$ GeV, and $0\leq\lambda_{2}\leq 8\pi$.
@@ 2.2 Two Higgs Doublet Models
We now considerthe second kind of Higgs sector extension of the Standard Model, namely the Two Higgs Doublet Models, in this paper. For a comprehensive review of the theory and phenomenological studies of the THDM, we refer the reader to Ref. [114] for further information. The model is briefly summarized in this section. A complex Higgs doublet with hypercharge $Y=1/2$ is added to the scalar sector of the SM. The scalar potential takes the form:
@@
[…]
Two scalar doublets are expanded around their VEVs after electroweak symmetry breaking (EWSB) as
@@
The vacuum expectation value is then fixed at $v=\sqrt{v_{1}^{2}+v_{2}^{2}}\sim 246$ GeV, in agreement with the SM value. After the EWSB, the physical particles in the THDM consist of two CP-even Higgs bosons, $h$ and $H$, with one of them, $h$, being identified as the SM-like Higgs boson discovered at the LHC, a CP-odd Higgs boson ($A^{0}$), and a pair of charged Higgs bosons ($H^{\pm}$). To obtain the physical masses of the new scalar bosons, we perform the following rotations:
@@
The mixing angle $\beta$ is defined by $t_{\beta}\equiv\tan\beta=v_{2}/v_{1}$. The physical masses of the Higgs bosons are then expressed in terms of the model parameters as follows:
@@
We present the relevant couplings involved in the amplitude computations for the processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in Tables 34 (for physical couplings) and in Table 5 (for unphysical couplings). These couplings are derived in Appendix B.